# vec2wav 2.0: Advancing Voice Conversion via Discrete Token Vocoders

### v3

@@ Abstract
We propose a new speech discrete token vocoder, vec2wav 2.0, which advances voice conversion (VC). We use discrete tokens from speech self-supervised models as the content features of source speech, and treat VC as a prompted vocoding task. To amend the loss of speaker timbre in the content tokens, vec2wav 2.0 utilizes the WavLM features to provide strong timbre-dependent information. A novel adaptive Snake activation function is proposed to better incorporate timbre into the waveform reconstruction process. In this way, vec2wav 2.0 learns to alter the speaker timbre appropriately given different reference prompts. Also, no supervised data is required for vec2wav 2.0 to be effectively trained. Experimental results demonstrate that vec2wav 2.0 outperforms all other baselines to a considerable margin in terms of audio quality and speaker similarity in any-to-any VC. Ablation studies verify the effects made by the proposed techniques. Moreover, vec2wav 2.0 achieves competitive cross-lingual VC even only trained on monolingual corpus. Thus, vec2wav 2.0 shows timbre can potentially be manipulated only by speech token vocoders, pushing the frontiers of VC and speech synthesis.
Index Terms:
@@
I Introduction
Discretizing speech into “tokens” has prevailed in speech generative tasks, such as text-to-speech (TTS) [1, 2, 3, 4], in the era of large language models (LLMs). However, the potential of discrete speech tokens in voice conversion (VC) has not been fully mined, which typically aims to convert source speech into target timbre from reference speech. Speech discrete tokens can be roughly divided into acoustic tokens and semantic tokens [5]. Although acoustic tokens [6, 7] reconstruct speech signals well, they lack the ability of VC because all aspects of information in speech are mixed and retained together. Semantic tokens usually come from speech self-supervised (SSL) models [8, 9, 10, 11] that emphasize on content-related information. Whether timbre is intentionally or unintentionally removed in these tokens, they can act as content representations and thus be utilized in the recognition-synthesis VC paradigm [12].
Throughout the history, VC methods with a continuous feature space have been researched with depth. The AutoVC series of work [13, 14, 15] attempt to decouple speech attributes via autoencoder bottlenecks. VC with advanced generative algorithms has then achieved remarkable results, such as normalizing flows [16, 17, 18] and diffusion models [19, 20, 21]. After the rise of speech SSL methods, researchers begin to apply SSL features in VC  [12, 22, 23, 24] where the rich phonetic content information from SSL features are utilized.
But continuous features make it hard to integrate VC with LLMs and thus an isolated step with other speech-related tasks. Discrete speech tokens can serve as content representations, so VC can be treated as a speech re-synthesis task then [25]. Starting from [25], discrete SSL features are increasingly investigated in VC for keeping phonetic content while discarding most of the acoustic details, such as UUVC [26] and Vec-Tok-VC+ [27]. ContentVec [28] introduces speaker disentanglement to SSL features that benefit VC. There also exist researches on decoupled speech codecs that also facilitate VC, such as SSVC [29] and FACodec [4]. Nevertheless, the performance of those VC methods is still limited compared to continuous state-of-the-arts. Also, excessive design of speaker disentanglement in the discrete tokens may lead to a negative impact on other paralinguistic information that needs to be preserved, such as prosody.
@@ I Introduction
  • Competitiveness. vec2wav 2.0 achieves superior any-to-any VC performance even compared to continuous VC methods. Furthermore, though only trained on English corpus, vec2wav 2.0 shows remarkable cross-lingual VC performance.
  • New Paradigm. vec2wav 2.0 proves that speaker timbre can be almost manipulated solely by vocoders even if the speech tokens are not perfectly speaker-decoupled, which may simplify the paradigm in the LLM-based zero-shot TTS world nowadays.
@@
II vec2wav 2.0: Prompted Token Vocoder
II-A System Overview
@@
After timbre is preliminarily merged into the frontend, we design an adaptive BigVGAN [31] generator to further incorporate the timbre embedding in waveform generation. The core component of this adaptive generator is a novel adaptive Snake activation function, which will be illustrated in Section II-B.
II-B Adaptive Snake Activation
@@
II-C Content and Prompt Features
Both the content and prompt inputs to vec2wav 2.0 are SSL features with different goals. For input discrete tokens, there should be as less timbre as possible; for prompt features, there should be enough and apparent timbre information to help reconstruction.
@@ II-C Content and Prompt Features
II-D Discriminators and Training Criterion
@@
The training criteria include the auxiliary mel prediction loss and all the other GAN losses from HifiGAN. The auxiliary mel prediction loss is an L1 loss between the ground truth mel-spectrograms and predicted ones that come from linear projections after the Conformer frontend, to warm up the whole model. This loss is added with a certain coefficient, and following vec2wav [1] and CTX-vec2wav [3], we cancel it after warming up.
@@ II-E Any-to-Any Voice Conversion
III Experiments
III-A Data and Model Setup
@@
The whole model is trained for 1 million steps on 4 NVIDIA A10 GPUs with a maximum batch capacity of 36s speech data per device. Other hyper-parameters follow CTX-vec2wav [3].
III-B English Any-to-Any VC
@@
To measure the performance of VC systems, we adopt multi-faceted metrics that contain both objective and subjective ones:
@@
    Quality and intelligibility: We use the subjective naturalness MOS (NMOS) and word error rate (WER percentage) between recognized hypotheses and ground truth texts. The NMOS tests require listeners to rate the utterances by quality and naturalness ranging from 1 to 5. WERs are computed using NeMo ASR [^4] https://huggingface.co/nvidia/stt_en_fastConformer_transducer_large.
@@
TABLE I: Performance comparison of English any-to-any VC
@@
vec2wav 2.0 4.51$\pm$0.09 3.29 4.46$\pm$0.08 0.886 0.722
We compare vec2wav 2.0 with some famous VC models. YourTTS [16] is a famous flow-based end-to-end VC model. DiffVC [19] and Diff-HierVC [20] promote convertibility via diffusion models. UUVC [26] also performs VC by discrete token reconstruction, but incorporates HuBERT tokens and additional prosody predictions. FACodec [4] is a state-of-the-art speech codec based on supervised decoupling of content, prosody, timbre and detail information. FACodec is capable of converting voices by simply replacing the speaker embedding into the target speaker and then decoding into waveform. We discard the detail tokens in FACodec for VC since we find these tokens still contain considerable speaker information that harms VC performance. We use the official implementation and checkpoints for all baselines. Note that the training data in all baselines either includes LibriTTS or is magnitudes larger (e.g. FACodec), so the comparisons are fair enough.
Table I presents the comparison results. “Source GT” means the ground truth source utterances, and MOS values are reported with 95% confidence intervals. It is thus clear that vec2wav 2.0 achieves significantly better synthesis quality and speaker similarity than all the baselines. The pitch correlation of vec2wav 2.0 is also at a high level666 Note that pitch correlation is less meaningful if speaker similarity is low.. The WER of vec2wav 2.0 is still acceptable although not the best. This is mostly due to the quantization errors that occur in the vq-wav2vec model itself.
III-C Cross-Lingual Any-to-Any VC
TABLE II: Performance comparison of cross-lingual any-to-any VC
@@ III-C Cross-Lingual Any-to-Any VC
vec2wav 2.0 4.47$\pm$0.07 3.39 4.33$\pm$0.07 0.846 0.684
To verify the cross-lingual VC ability of vec2wav 2.0, we use the same set of English source utterances in Table I, but convert to target speakers in other languages. We collect reference utterances from five languages777 Spanish, German, Dutch, Italian, French. in MLS [40]. The test set is the full combination of source and target utterances. For each of those languages, one male and one female speaker are randomly chosen as target speakers, and one reference utterance for each target speaker is sampled. We compare vec2wav 2.0 with the famous cross-lingual VC model YourTTS that is trained on multilingual data, and also Diff-HierVC which appears to be the most competitive in Table I. We conduct subjective and objective evaluations in the same way as Section III-B.
Table II shows the results. Although not trained on multilingual data, vec2wav 2.0 consistently outperforms YourTTS and Diff-HierVC in speaker similarity and quality with a significant margin. The WER and P.Corr comparisons show a similar conclusion with Table I that vec2wav 2.0 possesses a decent level of intelligibility and prosody preservation, although not the best. Therefore, it is demonstrated that vec2wav 2.0 performs competitive conversions, regardless of the languages of references.
III-D Ablation Study
Figure 3: Objective SECS and P.Corr comparisons with varied input tokens and models. Perfect VC systems should lie on the top right.
@@ III-D Ablation Study
To compare architectures, we additionally train two variants of vec2wav 2.0 on vq-wav2vec inputs: vec2wav 2.0-ab1 that replaces the adaptive Snake activations in BigVGAN by the original Snakes; and vec2wav 2.0-ab2 that further replaces BigVGAN with HifiGAN. Thus the comparison between vec2wav2.0 and “ab1” indicates the effect of adaptive Snake activation, while that between CTX-vec2wav and “ab2” shows the difference made by prompt feature and modules. We present the ablation studies in terms of SECS and P.Corr in Fig.3, together with the baselines in Section III-B. It can be found that vec2wav 2.0 obtains consistently large improvements in speaker similarity compared to the predecessor CTX-vec2wav in all the three input SSL tokens, while maintaining comparable pitch preservation. From the ablation of model architectures, it is obvious that the prompt-related improvements of vec2wav 2.0 make a substantial contribution to speaker similarity, while the adaptive Snake activations further advance the VC performance. The proposed vec2wav 2.0 with vq-wav2vec tokens is finally nearest to the top right corner of Fig.3, pushing the frontier of modern VC methods towards ideal voice converters.
III-E Analysis of Pitch Contours: Case Study
Pitch plays an important role in both the perception of global timbre and local prosody variations. An ideal VC system should retain pitch variations but shift the global ranges according to target speakers. Fig.4 shows the pitch contours of a sample converted by vec2wav 2.0. It can be seen that the average pitch of the converted utterance closely matches that of the reference, while the source and converted ones share the similar variations. This intuitively shows vec2wav 2.0 properly manages the global pitch range while keeping the local traits.
Figure 4: Pitch contours of a converted sample. Dashed lines refer to averages.
IV Conclusion
We present a novel VC method, vec2wav 2.0, based on the re-synthesis of speech discrete tokens. It takes advantage of SSL features in both content and timbre representations and enhances CTX-vec2wav in architectural designs. The adaptive Snake activation technique is proposed to better incorporate timbre into waveform reconstruction. The resulting model achieves remarkable performance on intra and cross-lingual VC tasks. We believe vec2wav 2.0 has promising impacts on the future LLM-based speech generation paradigm. Future efforts are needed in improving the intelligibility and prosody preservation of the proposed method.

### v4

@@ Abstract
We propose a new speech discrete token vocoder, vec2wav 2.0, which advances voice conversion (VC). We use discrete tokens from speech self-supervised models as the content features of source speech, and treat VC as a prompted vocoding task. To amend the loss of speaker timbre in the content tokens, vec2wav 2.0 utilizes the WavLM features to provide strong timbre-dependent information. A novel adaptive Snake activation function is proposed to better incorporate timbre into the waveform reconstruction process. In this way, vec2wav 2.0 learns to alter the speaker timbre appropriately given different reference prompts. Also, no supervised data is required for vec2wav 2.0 to be effectively trained. Experimental results demonstrate that vec2wav 2.0 outperforms all other baselines to a considerable margin in terms of audio quality and speaker similarity in English and cross-lingual any-to-any VC. Ablation studies verify the effects made by the proposed techniques.
keywords:
@@
1 Introduction
footnotetext: Kai Yu is the corresponding author.
Discretizing speech into “tokens” has prevailed in speech generative tasks, such as text-to-speech (TTS) [1, 2, 3, 4], in the era of large language models (LLMs). However, the potential of discrete speech tokens in voice conversion (VC) has not been fully mined, which typically aims to convert source speech into target timbre from reference speech. Speech discrete tokens can be roughly divided into acoustic tokens and semantic tokens [5]. Although general-purpose acoustic tokens [6, 7] reconstruct speech signals well, they lack the ability of VC because all aspects of information in speech are mixed and retained together. Semantic tokens usually come from speech self-supervised (SSL) models [8, 9, 10, 11] that emphasize on content-related information. No matter whether timbre is intentionally or unintentionally removed in these tokens, they can act as content representations and thus be utilized in the recognition-synthesis VC paradigm [12].
Among literature, VC methods with a continuous feature space have been researched with depth. These methods include speech decoupling via autoencoder bottlenecks [13, 14, 15], and the adoption of advanced generative algorithms like normalizing flow [16, 17] and diffusion models [18, 19, 20]. After the rise of speech SSL methods, researchers begin to apply SSL features in VC  [12, 21, 22, 23, 24, 25, 26] where the rich phonetic content information from SSL features are utilized.
However, VC with continuous features is hard to cooperate with LLMs, thus an isolated step from other speech-related tasks. Discrete speech tokens can also serve as content representations, thus VC can be treated as a speech re-synthesis task [27]. Recently, discrete SSL features are increasingly explored in VC to retain phonetic content while discarding most acoustic details [27, 28, 29]. There also exist researches on decoupling speech tokens that also facilitate VC, such as SSVC [30] and FACodec [4]. Nevertheless, the performance of those VC methods is still limited compared to continuous state-of-the-arts. Also, excessive design of speaker disentanglement in the discrete tokens may cause a negative impact on other paralinguistic information that needs to be preserved, such as prosody.
@@ 1 Introduction
  • Competitiveness. vec2wav 2.0 achieves superior any-to-any VC performance even compared to continuous VC methods and industry-level VC methods. Moreover, vec2wav 2.0 exhibits notable cross-lingual VC performance despite being trained only on English data.
  • New Paradigm. vec2wav 2.0 proves that speaker timbre can be almost manipulated solely by vocoders even if the speech tokens are not perfectly speaker-decoupled. This may simplify the paradigm of the LLM-based TTS world nowadays.
@@
2 vec2wav 2.0: Prompted Token Vocoder
2.1 System Overview
@@
After timbre is preliminarily merged into the frontend, we design an adaptive BigVGAN [32] generator to further incorporate the timbre embedding in waveform generation. The core component of this adaptive generator is a novel adaptive Snake activation function, which will be illustrated in Section 2.2.
2.2 Adaptive Snake Activation
@@
2.3 Content and Prompt Features
Both the content and prompt inputs to vec2wav 2.0 are SSL features with different goals: the input tokens should have as less timbre as possible, while the prompt features should contain sufficient and clear timbre information to aid reconstruction.
@@ 2.3 Content and Prompt Features
2.4 Discriminators and Training Criterion
@@
The training criteria include the auxiliary mel prediction loss and all the other GAN losses from HifiGAN. The auxiliary mel prediction loss is an L1 loss between the ground truth mel-spectrograms and predicted ones that come from linear projections after the Conformer frontend, to warm up the whole model. This loss is weighted with a certain coefficient, and we cancel it after warming up, following [1, 3].
@@ 2.5 Any-to-Any Voice Conversion
3 Experiments
3.1 Data and Model Setup
@@
The whole model is trained for 1 million steps on 4 NVIDIA A10 GPUs with a max batch size of 36s speech data per device. Other hyper-parameters follow CTX-vec2wav [3].
3.2 English Any-to-Any VC
@@
To comprehensively evaluate the performance of VC systems, we employ a range of objective and subjective metrics:
@@
    Quality and intelligibility: We use the subjective naturalness MOS (NMOS) and word error rate (WER) between ground truth and recognized texts. The NMOS tests require listeners to rate the utterances by quality and naturalness ranging from 1 to 5. WERs are computed using NeMo ASR [^4] https://huggingface.co/nvidia/stt_en_fastConformer_transducer_large.
@@
Table 1: Evaluation of English any-to-any VC
@@
[…]
Source GT 4.53$\pm$0.10 1.10 - - 1.000
Industry-level baseline
CosyVoice [41] 4.20$\pm$0.11 1.48 4.23$\pm$0.10 0.871 0.617
[…]
We compare vec2wav 2.0 with some famous VC models. YourTTS [16] is a famous flow-based end-to-end VC model. DiffVC [18] and Diff-HierVC [19] promote convertibility via diffusion models. UUVC [28] also performs VC by discrete token reconstruction, but incorporates HuBERT tokens and additional prosody predictions. FACodec [4] is a speech codec based on supervised decoupling of content, prosody, timbre and detail information. FACodec is capable of converting voices by simply replacing the speaker embedding into the target speaker and then decoding into waveform. We discard the detail tokens in FACodec for VC since we find these tokens still contain considerable speaker information that harms VC performance. We also conduct a separate listening test with a strong industry-level baseline, CosyVoice [41], which is trained on massive data. Its VC ability is based on the resynthesis of its supervised speech tokenizer using flow matching. We use the official checkpoints for all baselines. Note that the training data in all baselines either includes LibriTTS or is magnitudes larger (e.g. FACodec, CosyVoice), so the comparisons are fair enough.
Table 1 presents the comparison results. “Source GT” means source utterance recordings, and MOS values are reported with 95% confidence intervals. It is clear that vec2wav 2.0 achieves significantly higher synthesis quality and speaker similarity than all the academic baselines. Compared to CosyVoice, vec2wav 2.0 still owns slightly better performance in naturalness, similarity and prosody preservation, although the training data size is 300 times smaller. Its pitch correlation is also at a high level666 Note that pitch correlation is less meaningful if speaker similarity is low.. While the WER of vec2wav 2.0 is not the lowest, it remains acceptable. This is mostly due to the quantization errors inherent in the vq-wav2vec model itself.
3.3 Cross-Lingual Any-to-Any VC
Table 2: Evaluation of cross-lingual any-to-any VC
@@ 3.3 Cross-Lingual Any-to-Any VC
[…]
To verify the cross-lingual VC ability of vec2wav 2.0, we use the same set of English source utterances in Table 1, but convert to target speakers in other languages. We collect reference utterances from five languages777 Spanish, German, Dutch, Italian, French. in MLS [42]. The test set is the full combination of source and target utterances. For each of those languages, one male and one female speaker are randomly chosen as target speakers, and one reference utterance for each target speaker is sampled. We compare vec2wav 2.0 with the famous cross-lingual VC model YourTTS that is trained on multilingual data, and also Diff-HierVC which is a competitive academic baseline in Table 1. We conduct subjective and objective evaluations in the same way as Section 3.2.
Table 2 shows the results. Although not trained on multilingual data, vec2wav 2.0 consistently outperforms YourTTS and Diff-HierVC in speaker similarity and quality with a significant margin. The WER and P.Corr comparisons show a similar conclusion with Table 1 that vec2wav 2.0 possesses a decent level of intelligibility and prosody preservation, although not the best. Therefore, it is demonstrated that vec2wav 2.0 performs competitive conversions, regardless of the languages of references.
3.4 Ablation Study
Figure 3: Objective SECS and P.Corr comparisons with varied input tokens and models. Perfect VC systems should lie on the top right corner.
@@ 3.4 Ablation Study
To compare architectures, we additionally train two variants of vec2wav 2.0 on vq-wav2vec inputs: vec2wav 2.0-ab1 that replaces the adaptive Snake activations in BigVGAN by the original Snakes; and vec2wav 2.0-ab2 that further replaces BigVGAN with HifiGAN. Thus the comparison between vec2wav2.0 and “ab1” indicates the effect of adaptive Snake activation, while that between CTX-vec2wav and “ab2” shows the difference made by prompt feature and modules. We present the ablation studies in terms of SECS and P.Corr in Fig.3, together with the baselines in Section 3.2. It can be found that vec2wav 2.0 obtains consistently large improvements in speaker similarity compared to the predecessor CTX-vec2wav in all the three input SSL tokens, while maintaining comparable pitch preservation. From the ablation of model architectures, it is obvious that the prompt-related improvements of vec2wav 2.0 make a substantial contribution to speaker similarity, while the adaptive Snake activations further advance the VC performance. The proposed vec2wav 2.0 with vq-wav2vec tokens is finally nearest to the top right corner of Fig.3, pushing the frontier of modern VC methods towards ideal voice converters.
4 Conclusion
We present a novel VC method, vec2wav 2.0, based on the re-synthesis of speech discrete tokens. It takes advantage of SSL features in both content and timbre representations and enhances CTX-vec2wav in architectural designs. The adaptive Snake activation technique is proposed to better incorporate timbre into waveform reconstruction. The resulting model achieves remarkable performance on intra and cross-lingual VC tasks. We believe vec2wav 2.0 has promising impacts on the future LLM-based speech generation paradigm. Future efforts are needed in improving the intelligibility and prosody preservation of the proposed method, and the scaling ability on large-scale in-the-wild datasets needs to be explored.